# Weak expansion properties and a large deviation principle for coarse expanding conformal systems

### v2

@@ 6.1. Basic concepts
[…]
@@
A measurable partition $\mathcal{A}$ of $Z$ is a finite or countably infinite collection $\mathcal{A}=\{A_{j}:j\in J\}$ of mutually disjoint Borel sets with $\bigcup\mathcal{A}=Z$, where $J$ is a finite or countably infinite index set. For $x\in Z$ we denote by $\mathcal{A}(x)$ the unique element of $\mathcal{A}$ such that $x\in\mathcal{A}(x)$. Let $\mathcal{A}=\{A_{j}:j\in J\}$ and =$\mathcal{B}=\{B_{k}:k\in K\}$ be measurable partitions of $Z$, we say $\mathcal{A}$ is a refinement of $\mathcal{B}$ if for each $A_{j}\in\mathcal{A}$ there exist $B_{k}\in\mathcal{B}$ with $A_{j}\subseteq B_{k}$. The common refinement of $\mathcal{A}$ and $\mathcal{B}$ is defined as $\mathcal{A}\vee\mathcal{B}\coloneqq\{A_{j}\cap B_{k}:A_{j}\in\mathcal{A},\ B_{k}\in\mathcal{B}\}$, which is also a measurable partition. For a continuous map $g\colon Z\rightarrow Z$, we set $g^{-1}(\mathcal{A})\coloneqq\{g^{-1}(A_{j}):A_{j}\in\mathcal{A}\}$, and denote for $n\in\mathbb{N}$,
@@
$H_{\mu}(\mathcal{A})\coloneqq\sum_{j\in J}-\mu(Aj)\log(\mu(A_{j}))$
where $0\log 0$ is defined to be 00. If $H_{\mu}(\mathcal{A})<+\infty$, then the limit $h_{\mu}(g,\mathcal{A})\coloneqq\lim\limits_{n\to+\infty}\frac{1}{n}H_{\mu}(\mathcal{A}_{g}^{n})$ exists, and the measure-theoretic entropy of $g$ for $\mu$ is given by
@@
In particular, when $\psi$ is the constant function 00, we have
@@ 6.2. Existence of equilibrium states
Let $Z$ be a compact metric space and $g\colon Z\rightarrow Z$ a continuous map. First, we recall that $\mathcal{M}(Z,g)$ is $*{}^{*}$ compact. We refer the reader to [PU10] for more details.
@@
is upper semi-continuous with respect to the $*{}^{*}$ topology.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{0}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4, and $\psi\in C(X)$. Note that $X$ is compact. Denote $f_{X}\coloneqq f|_{X}$.
By Proposition 6.1 the map $h_{\centerdot}(f_{X})\colon\mathcal{M}(X,f_{X})\rightarrow\mathbb{R},\ \mu\mapsto h_{\mu}(f_{X})$ is upper semi-continuous. Then since $\psi\in C(X)$, the map
$P_{\centerdot}(f_{X},\psi)\colon\mathcal{M}(X,f_{X})\rightarrow\mathbb{R},\ \mu\mapsto P_{\mu}(X,f_{X})=h_{\mu}(f_{X})+\int\!\psi{\,\mathrm{d}}\mu$
is upper semi-continuous. Since $\mathcal{M}(X,f_{X})$ is $*{}^{*}$ compact, there exists $\nu\in\mathcal{M}(X,f_{X})$ that attains the supremum in $\sup\{P_{\mu}(f_{X},\psi):\mu\in\mathcal{M}(X,f_{X})\}$. By the Variational Principle, such $\nu\in\mathcal{M}(X,f_{X})$ is an equilibrium state. ∎
In particular, let $\psi$ be the constant function 00, then we obtain the existence of measures of maximal entropy.
@@ 7.1. Level-$2$ large deviation principle
Let $Z$ be a compact metric space, and $\mathcal{P}(Z)$ be the set of Borel probability measures on $Z$ equipped with the $*{}^{*}$ topology. Note this topology is metrizable (see for example, [Co85, Theorem 5.1]). Let $I\colon\mathcal{P}(Z)\rightarrow[0,+\infty]$ be a lower semi-continuous function.
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log(\Omega_{n}(\mathfrak{F}))$ $\leq-\inf\{I(x)\colon x\in\mathfrak{F}\},$
$\liminf_{n\to+\infty}\frac{1}{n}\log(\Omega_{n}(\mathfrak{G}))$ $\geq-\inf\{I(x)\colon x\in\mathfrak{G}\}.$
@@
Then $\{\Omega_{n}\}_{n\in\mathbb{N}}$ satisfies a level-$2$ large deviation principle with rate function $I^{\phi}$, and it converges in the $*{}^{*}$ topology to the Dirac measure supported on the unique equilibrium state for $g$ and potential $\phi$. Furthermore, for each convex open subset $\mathfrak{G}$ of $\mathcal{P}(Z)$ containing some invariant measure, we have
@@ 7.2. The Additional Assumptions
Before we state these assumptions, we recall the notion of strongly path-connectedness. See [DPTUZ21, Definition 2.3].
@@
These additional assumptions were adopted in [DPTUZ21].
@@ 7.3. Proof of Theorem D
The goal of this subsection is to establish Theorem D. We first recall the notion of geometric coding tree from [DPTUZ21] and then use it to establish some characterizations of the topological pressure before proving Theorem D at the end of this subsection.
In the following context, we assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2, and let $\phi$ be a real-valued Hölder continuous function on $X$. We denote $f_{X}\coloneqq f|_{X}$ for convenience.
@@
We first verify that for each $\psi\in H$,
@@
By results in [DPTUZ21] that, under the Additional Assumption in Subsection 7.2, the topological pressure of a metric CXC system and the shift map on a space of one-sided sequences can be related through a semiconjugacy. We will use this idea to verify (7.1).
@@ 7.3.1. Geometric coding tree and semiconjugacy
Let $\Sigma\coloneqq\{1,\,\dots,\,d\}^{\mathbb{N}}$ be the space of infinite one-sided sequences on $d$ symbols, and $\sigma\colon\Sigma\rightarrow\Sigma$ the shift map. We equip $\Sigma$ with the standard metric $\rho(\xi,\eta)\coloneqq 2^{-\min\{k\geq 0\,:\,\xi_{k}\neq\eta_{k}\}}$ for distinct $\xi=(\xi_{0},\xi_{1},\dots,)$ and $\eta=(\eta_{0},\eta_{1},\dots)$ in $\Sigma$.
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Pick $w\in X\setminus\operatorname{post}(f)$, let $w_{1},w_{2},\dots,w_{d}$ be all of its preimages under $f$. For each $w_{i}$, choose a path $\gamma_{i}$ in $\mathfrak{X}_{0}$ connecting $w$ to $w_{i}$ and avoiding $\operatorname{post}(f)$. Such paths exist since $\mathfrak{X}_{1}$ and $\mathfrak{X}_{0}$ are strongly path-connected. Here, by a path $\gamma$ in $\mathfrak{X}_{0}$ we mean a continuous map $\gamma\colon[0,1]\rightarrow\mathfrak{X}_{0}$.
@@
Actually, such lifts are unique, which is well-known to experts. We include a proof for the convenience of the reader.
@@
[…]
Proof.
Consider $J\coloneqq\{t\in[0,1]:\widetilde{\gamma}_{1}(t)=\widetilde{\gamma}_{2}(t)\}$. Then $0\in J$. It suffices to prove that $J$ is an open and closed subset of $[0,1]$.
[…]
Fix an arbitrary sequence $\{t_{n}\}_{n\in\mathbb{N}}$ of points in $J$ such that $\lim_{n\to+\infty}t_{n}=t^{*}\in[0,1]$. Then $\widetilde{\gamma}_{1}(t_{n})=\widetilde{\gamma}_{2}(t_{n})$ for each $n\in\mathbb{N}$. Since $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ are continuous, we have $\widetilde{\gamma}_{1}(t^{*})=\widetilde{\gamma}_{2}(t^{*})$. Thus $t^{*}\in J$. Since the sequence $\{t_{n}\}_{n\in\mathbb{N}}$ is arbitrary, $J$ is closed.
Since $J$ is open and closed, and $J\neq\emptyset$, we have $J=[0,1]$. Now the proof is complete. ∎
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Let $\xi\in\Sigma$. Then for each pair of $m,\,k\in\mathbb{N}$ with $k\leq m$, we have $f^{k}(z_{m}(\xi))=z_{m-k}\bigl{(}\sigma^{k}(\xi)\bigr{)}$.
@@
By induction, $f(z_{m}(\xi))=z_{m-1}(\sigma(\xi))$ holds for all $m\in\mathbb{N}$. Then it follows immediately that $f^{k}(z_{m}(\xi))=z_{m-k}\bigl{(}\sigma^{k}(\xi)\bigr{)}$ holds for each pair of $m,\,k\in\mathbb{N}$ with $k\leq m$, and the proof is complete. ∎
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Let $\xi=(\xi_{0},\,\xi_{1},\,\cdots)$ and $\eta=(\eta_{0},\,\eta_{1},\,\cdots)$ be in $\Sigma$. If $\xi_{n}\neq\eta_{n}$ for some $n\in\mathbb{N}_{0}$, then $z_{m}(\xi)\neq z_{m}(\eta)$ for each $m\in\mathbb{N}$ with $m\geq n$.
@@
Assume that $\xi_{n}\neq\eta_{n}$ for some $n\in\mathbb{N}_{0}$. We prove that $z_{m}(\xi)\neq z_{m}(\eta)$ for all $m\geq n$ by induction.
@@
[…]
By induction, $z_{m}(\xi)\neq z_{m}(\eta)$ holds for all $m\geq n$. ∎
@@
Fix arbitrary $x_{0}\in X\setminus\operatorname{post}(f)$, $\omega_{0}\in\Sigma$, $n\in\mathbb{N}$, and $\xi,\eta\in\sigma^{-n}(\omega_{0})$ with $\xi\neq\eta$ and $x_{0}=\pi(\omega_{0})$ (see Proposition 7.6). By Lemma 7.9, for each integer $m\geq n$, it holds that $f^{n}(z_{m}(\xi))=z_{m-n}(\omega_{0})=f^{n}(z_{m}(\eta))$. On the other hand, since $\xi\neq\eta$, there exists $k\in\{0,1,\cdots,n-1\}$ such that $\xi_{k}\neq\eta_{k}$. Then by Lemma 7.10, we know that $z_{m}(\xi)\neq z_{m}(\eta)$ holds for each integer $m\geq n$.
@@
Thus we know that $\deg(f^{n};x)\geq 2$, which contradicts to $x\in{X\setminus\operatorname{post}(f)}$. Since $\xi,\,\eta$ are arbitrary, the map $\pi|_{\sigma^{-n}(\omega_{0})}\colon\sigma^{-n}(\omega_{0})\rightarrow f^{-n}(x_{0})$ is injective. Since $x_{0}\in X\setminus\operatorname{post}(f)$, we have $\operatorname{card}(f^{-n}(x_{0}))=d^{n}=\operatorname{card}(\sigma^{-n}(\omega_{0}))$, and we get that the map $\pi|_{\sigma^{-n}(\omega_{0})}$ is bijective. ∎
@@
$\frac{1}{C_{0}}\leq\frac{\sum_{\xi\in\sigma^{-n}(\xi_{0})}\exp(S_{n}\phi(\xi))}{\sum_{\eta\in\sigma^{-n}(\eta_{0})}\exp(S_{n}\phi(\eta))}\leq C_{0}.$

### v3

@@ 6.1. Basic concepts
$d_{g}^{n}(x,y)\coloneqq\max\bigl{\{}d\bigl{(}g^{k}(x),g^{k}(y)\bigr{)}:k\in\{0,\,1,\,\dots,\,n-1\}\bigr{\}},$
@@
A measurable partition $\mathcal{A}$ of $Z$ is a finite or countably infinite collection $\mathcal{A}=\{A_{j}:j\in J\}$ of mutually disjoint Borel sets with $\bigcup\mathcal{A}=Z$, where $J$ is a finite or countably infinite index set. For $x\in Z$ we denote by $\mathcal{A}(x)$ the unique element of $\mathcal{A}$ such that $x\in\mathcal{A}(x)$. Let $\mathcal{A}=\{A_{j}:j\in J\}$ and $\mathcal{B}=\{B_{k}:k\in K\}$ be measurable partitions of $Z$, we say $\mathcal{A}$ is a refinement of $\mathcal{B}$ if for each $A_{j}\in\mathcal{A}$ there exist $B_{k}\in\mathcal{B}$ with $A_{j}\subseteq B_{k}$. The common refinement of $\mathcal{A}$ and $\mathcal{B}$ is defined as $\mathcal{A}\vee\mathcal{B}\coloneqq\{A_{j}\cap B_{k}:A_{j}\in\mathcal{A},\ B_{k}\in\mathcal{B}\}$, which is also a measurable partition. For a continuous map $g\colon Z\rightarrow Z$, we set $g^{-1}(\mathcal{A})\coloneqq\{g^{-1}(A_{j}):A_{j}\in\mathcal{A}\}$, and denote for $n\in\mathbb{N}$,
@@
$H_{\mu}(\mathcal{A})\coloneqq\sum_{j\in J}-\mu(A_{j})\log(\mu(A_{j}))$
where $0\log 0$ is defined to be 0. If $H_{\mu}(\mathcal{A})<+\infty$, then the limit $h_{\mu}(g,\mathcal{A})\coloneqq\lim\limits_{n\to+\infty}\frac{1}{n}H_{\mu}\bigl{(}\mathcal{A}_{g}^{n}\bigr{)}$ exists, and the measure-theoretic entropy of $g$ for $\mu$ is given by
@@
In particular, when $\psi$ is the constant function 0, we have
@@ 6.2. Existence of equilibrium states
Let $Z$ be a compact metric space and $g\colon Z\rightarrow Z$ a continuous map. First, we recall that $\mathcal{M}(Z,g)$ is weak compact. We refer the reader to [PU10] for more details.
@@
is upper semi-continuous with respect to the weak topology.
@@
Let $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{0}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4, and $\phi\in C(X)$. Note that $X$ is compact.
By Proposition 6.1 and Theorem A, the map $h_{\centerdot}(f_{X})\colon\mathcal{M}(X,f_{X})\rightarrow\mathbb{R},\ \mu\mapsto h_{\mu}(f_{X})$ is upper semi-continuous. Then since $\phi\in C(X)$, the map
$P_{\centerdot}(f_{X},\phi)\colon\mathcal{M}(X,f_{X})\rightarrow\mathbb{R},\ \mu\mapsto P_{\mu}(X,f_{X})=h_{\mu}(f_{X})+\int\!\phi{\,\mathrm{d}}\mu$
is upper semi-continuous. Since $\mathcal{M}(X,f_{X})$ is weak compact, there exists $\nu\in\mathcal{M}(X,f_{X})$ that attains the supremum in $\sup\{P_{\mu}(f_{X},\phi):\mu\in\mathcal{M}(X,f_{X})\}$. By the Variational Principle, such $\nu\in\mathcal{M}(X,f_{X})$ is an equilibrium state. ∎
In particular, let $\phi$ be the constant function 0, then we obtain the existence of measures of maximal entropy.
@@ 7.1. Level-$2$ large deviation principle
Let $Z$ be a compact metric space, and $\mathcal{P}(Z)$ be the set of Borel probability measures on $Z$ equipped with the weak topology. Note this topology is metrizable (see for example, [Co85, Theorem 5.1]). Let $I\colon\mathcal{P}(Z)\rightarrow[0,+\infty]$ be a lower semi-continuous function.
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log(\Omega_{n}(\mathfrak{F}))$ $\leqslant-\inf\{I(x)\colon x\in\mathfrak{F}\},$
$\liminf_{n\to+\infty}\frac{1}{n}\log(\Omega_{n}(\mathfrak{G}))$ $\geqslant-\inf\{I(x)\colon x\in\mathfrak{G}\}.$
@@
Then $\{\Omega_{n}\}_{n\in\mathbb{N}}$ satisfies a level-$2$ large deviation principle with rate function $I^{\phi}$, and it converges in the weak topology to the Dirac measure supported on the unique equilibrium state for $g$ and potential $\phi$. Furthermore, for each convex open subset $\mathfrak{G}$ of $\mathcal{P}(Z)$ containing some invariant measure, we have
@@ 7.2. The Additional Assumptions
Before we state these assumptions, we recall the notion of strongly path-connectedness. Here we adopt [DPTUZ21, Definition 2.3].
@@
Note that these additional assumptions were adopted in [DPTUZ21].
@@ 7.3. Proof of Theorem D
The goal of this subsection is to establish Theorem D. We first recall the notion of geometric coding tree from [DPTUZ21] and then use it to establish a characterization of the topological pressure. With such a characterization, we can apply Theorem 7.1 to prove Theorem D.
In the following context, we assume that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2, and let $\phi$ be a real-valued Hölder continuous function on $X$.
@@
We first verify that for each $\psi\in C^{0,\alpha}(X)$,
@@
By results in [DPTUZ21], under the Additional Assumption in Subsection 7.2, the topological pressure of a metric CXC system and the shift map on a space of one-sided sequences can be related through a semiconjugacy. We will use this idea to verify (7.1).
@@ 7.3.1. Geometric coding tree and semiconjugacy
Let $\Sigma\coloneqq\{1,\,\dots,\,d\}^{\mathbb{N}}$ be the space of infinite one-sided sequences on $d$ symbols, and $\sigma\colon\Sigma\rightarrow\Sigma$ the shift map. We equip $\Sigma$ with the standard metric $\rho(\xi,\eta)\coloneqq 2^{-\min\{k\geqslant 0\,:\,\xi_{k}\neq\eta_{k}\}}$ for distinct $\xi=(\xi_{0},\xi_{1},\dots,)$ and $\eta=(\eta_{0},\eta_{1},\dots)$ in $\Sigma$.
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Pick $w\in X\setminus\operatorname{post}(f)$, let $w_{1},w_{2},\dots,w_{d}$ be all of its preimages under $f$. For each $w_{i}$, choose a path $\gamma_{i}$ in $\mathfrak{X}_{0}$ connecting $w$ to $w_{i}$ and avoiding $\operatorname{post}(f)$. Such paths exist since $\mathfrak{X}_{0}$ is strongly path-connected. Here, by a path $\gamma$ in $\mathfrak{X}_{0}$ we mean a continuous map $\gamma\colon[0,1]\rightarrow\mathfrak{X}_{0}$.
@@
Actually, such a lift is unique. We skip the proof since it is standard.
@@
Let $h\colon Y\rightarrow Z$ be a finite branched covering map, and $\gamma$ be a continuous path in $Z$ that is disjoint from the set of branch values $V_{h}$. Let $x=\gamma(0)$, and $\widetilde{x}\in h^{-1}(x)$. Let $\widetilde{\gamma}_{1}$ and $\widetilde{\gamma}_{2}$ be two paths in $Y$ such that $h(\widetilde{\gamma}_{i})=\gamma$ and $\widetilde{\gamma}_{i}(0)=\widetilde{x}$, $i\in\{1,\,2\}$, then $\widetilde{\gamma}_{1}=\widetilde{\gamma}_{2}$.
@@
It can be checked that when $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2, the conditions in [DPTUZ21, Theorem 1.1, Proposition 2.16, and Lemma 4.2] are satisfied. Thus, we obtain what follows.
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Let $\xi=(\xi_{0},\,\xi_{1},\,\cdots)$ be in $\Sigma$. Then for each pair of $m,\,k\in\mathbb{N}$ with $k\leqslant m$, we have $f^{k}(z_{m}(\xi))=z_{m-k}\bigl{(}\sigma^{k}(\xi)\bigr{)}$.
@@
By induction, $f(z_{m}(\xi))=z_{m-1}(\sigma(\xi))$ holds for all $m\in\mathbb{N}$. Then it follows immediately that $f^{k}(z_{m}(\xi))=z_{m-k}\bigl{(}\sigma^{k}(\xi)\bigr{)}$ holds for each pair of $m,\,k\in\mathbb{N}$ with $k\leqslant m$, and the proof is complete. ∎
@@
Suppose that $f\colon(\mathfrak{X}_{1},X)\rightarrow(\mathfrak{X}_{0},X)$ and $\{\mathcal{U}_{n}\}_{n\in\mathbb{N}_{0}}$ satisfy the Assumptions in Section 4 and the Additional Assumptions in Subsection 7.2. Let $\xi=(\xi_{0},\,\xi_{1},\,\cdots)$ and $\eta=(\eta_{0},\,\eta_{1},\,\cdots)$ be in $\Sigma$. If $\xi_{n}\neq\eta_{n}$ for some $n\in\mathbb{N}_{0}$, then $z_{m}(\xi)\neq z_{m}(\eta)$ for each $m\in\mathbb{N}$ with $m\geqslant n$.
@@
Assume that $\xi_{n}\neq\eta_{n}$ for some $n\in\mathbb{N}_{0}$. We prove that $z_{m}(\xi)\neq z_{m}(\eta)$ for all $m\geqslant n$ by induction.
@@
For an integer $m>n$, we make the induction hypothesis that $z_{m-1}(\xi)\neq z_{m-1}(\eta)$. Suppose $z_{m}(\xi)=z_{m}(\eta)$. Then $w_{i_{m}}=f^{m}(z_{m}(\xi))=f^{m}(z_{m}(\eta))$ for some $i_{m}\in\{1,\,2,\,\dots,\,d\}$. Then $\gamma_{m}(\xi)$ and $\gamma_{m}(\eta)$ are two paths such that $f^{m}(\gamma_{m}(\xi))=f^{m}(\gamma_{m}(\eta))=\gamma_{i_{m}}$ and $\gamma_{m}(\xi)(1)=z_{m}(\xi)=z_{m}(\eta)=\gamma_{m}(\eta)(1)$. By Lemma 7.5, $\gamma_{m}(\xi)=\gamma_{m}(\eta)$. Thus $z_{m-1}(\xi)=z_{m-1}(\eta)$, which contradicts to our induction hypothesis. So $z_{m}(\xi)\neq z_{m}(\eta)$.
By induction, $z_{m}(\xi)\neq z_{m}(\eta)$ holds for all $m\geqslant n$. ∎
@@
Fix arbitrary $x_{0}\in X\setminus\operatorname{post}(f)$, $\omega_{0}\in\Sigma$, $n\in\mathbb{N}$, and $\xi,\eta\in\sigma^{-n}(\omega_{0})$ with $\xi\neq\eta$ and $x_{0}=\pi(\omega_{0})$ (see Proposition 7.6). By Lemma 7.9, for each integer $m\geqslant n$, it holds that $f^{n}(z_{m}(\xi))=z_{m-n}(\omega_{0})=f^{n}(z_{m}(\eta))$. On the other hand, since $\xi\neq\eta$, there exists $k\in\{0,1,\cdots,n-1\}$ such that $\xi_{k}\neq\eta_{k}$. Then by Lemma 7.10, we know that $z_{m}(\xi)\neq z_{m}(\eta)$ holds for each integer $m\geqslant n$.
@@
Thus we know that $\deg(f^{n};x)\geqslant 2$, which contradicts to $x_{0}\in{X\setminus\operatorname{post}(f)}$. Since $\xi,\,\eta$ are arbitrary, the map $\pi|_{\sigma^{-n}(\omega_{0})}\colon\sigma^{-n}(\omega_{0})\rightarrow f^{-n}(x_{0})$ is injective. Since $x_{0}\in X\setminus\operatorname{post}(f)$, we have $\operatorname{card}(f^{-n}(x_{0}))=d^{n}=\operatorname{card}(\sigma^{-n}(\omega_{0}))$, and we get that the map $\pi|_{\sigma^{-n}(\omega_{0})}$ is bijective. ∎
@@
$\frac{1}{C_{0}}\leqslant\frac{\sum_{\xi\in\sigma^{-n}(\xi_{0})}\exp(S_{n}\phi(\xi))}{\sum_{\eta\in\sigma^{-n}(\eta_{0})}\exp(S_{n}\phi(\eta))}\leqslant C_{0}.$